OpenStudy (anonymous):

Solve the following quadratic equation using the quadratic formula and then choose the correct solution set. 6x2 - 7x + 2 = 0

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

@dominique2525 Thats not one of the options. These are the options: {-1/2,2/3} {1/2,2/3} {1/2,3/2}

OpenStudy (anonymous):

I got dominiques answer too, but how come its not one of the choices? :/

OpenStudy (campbell_st):

I think there is a significant error in the calculations \[x =\frac{ 7 \pm \sqrt{(-7)^2 - 4 \times 6 \times 2}}{2 \times 6}\] which becomes \[x = \frac{7 \pm \sqrt{49 - 48}}{12}\]

OpenStudy (anonymous):

How do i figure out what the solution set is?

OpenStudy (campbell_st):

and just to show your teacher, this quadratic can be factored multipily a and c 6 x 2 = 12 find the factors of 12 that add to -7.... they are both negative -3, -4 then write the factored equation as (6x -3)(6x -4) ------------ 6 factor the linear factors 3(2x -1) 2(3x -2) --------------- 6 remove the common factors (2x -1)(3x -2) = 0 now you can find the values

OpenStudy (campbell_st):

well to get the solution simplify the square root \[x = \frac{7 + \sqrt{1}}{12}~~~~and~~~~~ x = \frac{7 - \sqrt{1}}{12}\] what is the value of \[\sqrt{1} = ?\]

OpenStudy (anonymous):

1?

OpenStudy (campbell_st):

great so you have 2 answers \[x = \frac{7 + 1}{12} ~~~~and~~~~x = \frac{7 -1}{12}\] just simplify them

OpenStudy (anonymous):

@campbell_st Thank you so much!!! ugh i finally understand!!!!!

OpenStudy (campbell_st):

you made the calculation error inside the square root (-7)^2 - 4 x 6 x 2 you someone posted -49 - 48 = 97 it should 49 - 48 = 1